Write down the equation for the line that passes through the point (3, -2) and is perpendicular to the line y = 2x + 4.
How do I do this?

Answer by suresh(20)

To find the equation of line passing through (3,-2)
and perpendicular to y=2x+4
We know that the perpendicular line for ax+by+c = 0 is -bx+ay+k=0
So the perpendicular line to 2x-y+4=0 is x+2y+k =0 ----->(1)
We have to find the value of k.
Since the line passing through the point(3,-2),it satisfies the equation.
plug x=3 and y = -2 in (1)
3+2(-2)+k =0.
-1 + k = 0.
k=1.
Plug the value of k in (1) we get the required line.
Therefore the required line is, x+2y+1=0.